OpenStudy (anonymous):

Solve the equation for S 3s^2+22s+40=0

OpenStudy (anonymous):

\[(s+4) (3 s+10) = 0\] is a start

OpenStudy (anonymous):

3s^2+12s+10s+40=0 3s(s+4) +10(s+4)=0 (3s+10)(s+4)=0 s=-10/3 or s=-4